The ratio of the price of coffee A to coffee B is 2:3. During a certain month, a store sells coffee A and coffee B in the ratio of 4:5. If the total revenue from the sale of these two coffees is $27,600, how much of the revenue was from the sale of coffee A?
Please explain.
Thanks.

Answer by mahalingk(2)

Price ratio of A,B is 2:3 in general m:n
sold out ratio of A,B is 4:5 in general p:q
revenue ratio of A,B always will be 2*4:3*5 in general mp:nq
As total revenue is 27000, and their ratio is 8:15 say(8x,15x)
we have the exact revenue as follows, 8x+15x=27600
23x=27600
x=27600/23 ==> x=1200 that is we have the exact revenue ratio as of A is 8x=8*1200=9600
B is 15x=15*1200=18000